At what minimum speed must a roller coaster be traveling when upside down at the top of a circle if the passengers are not to fall out. The radius of curvature is 5.6 metres. The mass of the passenger is 72.00 kg. [Answer in km/h]
Answer #1

minimum safe velocity

v^2 = rg

v^2 = 5.6*9.8

v = 7.408 m/s

v = 7.408* 18 / 5

v = 26.67 km/h

======

Comment in case any doubt, will reply for sure . Goodluck